OpenStudy (anonymous):

What is the slope of the line 4x + 6y = 12?

OpenStudy (anonymous):

So if you switch it from standard form to slope-intercept you; 1) isolate the 'y' intercept 6y = -4x + 12 2) divide each side by 6 y = (-4x + 12) / (6) 3) determine the solution y = -0.67x + 2

OpenStudy (anonymous):

\[4x+6y=12 => 6y=-4x+12 => y=-\frac{ 4 }{ 6 }x+2\] compare it with standard eq. (y=mx+c) so slope is -4/6

OpenStudy (anonymous):

remember what I told you m = slope so what's the slope?

OpenStudy (anonymous):

4x? i suck at this!

OpenStudy (anonymous):

Lol you're doing fine, the slope-intercept form is y = mx + b where 'm' is the slope what number came before 'x' in the equation?

OpenStudy (anonymous):

these are the choices A. 2/3 B. 2 C. -2/3 D. 6 i dont know which one to pick ugh >.<

OpenStudy (anonymous):

-2/3 is equal to -0.67 AND -4/6

OpenStudy (anonymous):

ohh yeah i get it thanks

OpenStudy (anonymous):

its -2/3..

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

took me a while,

OpenStudy (anonymous):

It took me a loooong while to get it down, but after a few years of algebra algebra 2 and geometry, it's easier >.<

OpenStudy (anonymous):

this is my 2nd year i'm still pretty lost with slope and this kind of stuff haha

OpenStudy (anonymous):

Don't fret, you'll get it :)

OpenStudy (anonymous):

i'm going to post 2 more and that will be it :)

OpenStudy (anonymous):

Okay :)